Solve the problem below. Show your solution.
A father is now 9 times as old as his son. In 3 years, the father’s age will be 5 times the son’s age.
How old is the son now?

Answer by ikleyn(52788)

.

Father  9x

Son      x



In 3 years

    9x + 3 = 5*(x+3)


Simplify and solve

    9x + 3 = 5x + 15

    9x - 5x = 15 - 3

      4x    =    12

       x    =    12/4 = 3.


ANSWER.  The son is 3 years old.  The father is 9*3 = 27 years old.

Solved, explained, answered and completed.